In printing an article of 48,000 words, a printer decides to use two sizes of type. Using the larger type, a printed page contains 1,800 words. Using a smaller type, a page contains 2,400 words. The article is allotted 21 full pages in a magazine. How many pages must be
in smaller type? ___________

Answer by checkley77(12844)

2400X+1800(21-X)=48000
2400X+37,800-1800X=48,000
600X=48,000-37,800
600X=10,200
X=10,200/600
X=17 PAGES OF SMALL PRINT.
21-17=4 PAGES OF LARGE PRINT.
PROOF:
2400*17+1800*4*=48,000
40,800+7,200=48,000
48,000=48,000